twice a number plus three times the sum of 2 and the number is thirty one. what is the number/
a number squared is 625. what is that number

Answer by Greenfinch(383)

2x +3(x+2) = 31
5x + 6 = 31
5x = 25
x = 5
x^2 = 625 = 5^4
x = 25